OpenStudy (anonymous):

May you please help me with this problem? @Loser66

OpenStudy (yanasidlinskiy):

(-5,-2)

OpenStudy (anonymous):

How? @YanaSidlinskiy

OpenStudy (yanasidlinskiy):

*I think* Don't take my word for it, though. You again probably want to ask, @undeadknight26

undeadknight26 (undeadknight26):

We are scaling it down by three seeing that: 18/3 = 6 (counting blocks) 9/3 = 3 6/3 = 2 Giving us the points of: (-4,-5) Im not sure da banana scaled it down though!

OpenStudy (anonymous):

^Oh okay and thank you for tagging someone.

OpenStudy (anonymous):

da banana?

undeadknight26 (undeadknight26):

@YanaSidlinskiy lol sorry XD

undeadknight26 (undeadknight26):

I call yana the banana.

OpenStudy (anonymous):

Aww. -So you're sure it's \[(-4,-5) \] and not \[ (-5,-2) \] ?

undeadknight26 (undeadknight26):

Not 100% sure...Not sure what @YanaSidlinskiy did to get her coordinates but most likely it was something smart...I would say its a 50/50 chance on our answers.

OpenStudy (anonymous):

:(

undeadknight26 (undeadknight26):

Use your instincts oh superior one! Did they not teach you this in your lesson? What am i talkin bout i dun even read my lesson -_-

OpenStudy (anonymous):

I'm not superior.. I don't exactly understand it. Most of the time I just have someone check my answers, but I too didn't understand this one. -Lmbo.

OpenStudy (yanasidlinskiy):

Ughghghhgh!!!! I'm out! Why does everyone cal bannana! leaving for life!!!! neva coming back! Don't expect me 2 come bck either. Counted downxD! @undeadknight26 That's wut i did!!

undeadknight26 (undeadknight26):

Only i may call you banana -_-